# Towards General Conceptual Model Editing via Adversarial Representation Engineering

### v1

@@ 19       $I\leftarrow I_{T}\cup I_{A};$
To provide a clear understanding of the alternative training process, we offer a visualization in Figure 3. We compiled a set of 256 prompts, evenly divided between normal and malicious, with the expectation that the aligned model will reject all malicious inputs. The representations derived from these prompts are plotted using t-SNE, as depicted in the figure.
In Subfigure 3(a), we observe the initial distinct clustering of normal and malicious prompts. Our goal for model editing is to adjust these representations so that the two types of prompts yield similar responses. During the first epoch, illustrated in Subfigure 3(b), the malicious prompts begin to converge towards the cluster of normal prompts. Since the two categories of representations remain distinct, necessitating further refinement of the discriminator. After 30 epochs of iterative training as shown in Subfigure 3(c), we observe that the representation of normal prompts remains consistent, having been continuously classified correctly. Meanwhile, the representations of malicious prompts have nearly merged into the normal cluster, making it challenging for the classifier to distinguish them. At this stage, the differences in representations are minimal and can be considered negligible, indicating a successful editing process.
@@ 4.2 General Conceptual Editing
Step 1. Train $D(\delta)$ to fit $G$ by optimizing $\mathcal{L}(\delta)$: Consider generated target representations $R_{T}$ corresponding to $c^{+}$ and antithetical representations corresponding to $c^{-}$. The loss $\mathcal{L}(\delta)$ is defined as the classic cross-entropy loss, which is
$\mathcal{L}(\delta)=\frac{1}{|R_{T}\cup R_{A}|}(\sum_{r\in R_{T}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{+}]+\sum_{r\in R_{A}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{-}]);$
@@
$\mathcal{L}(\theta)=\frac{1}{|I|}(\sum_{x\in I}-\log\mathbb{P}_{\theta}[D(G_{\theta}(x))=c^{+}]).$
Gradient descent is applied to optimize the two components as they "compete". Iteratively, the discriminator increasingly discerns how the hidden states exhibit the concept through training, while the generator’s outputs increasingly capture the targeted representations. Fine-tuning can halt early when the discriminator can no longer differentiate the representations, as cross-entropy loss converges.
@@ 5.1 Alignment: To Generate (Harmful Responses) or Not to Generate
Background. With the application of various alignment techniques [27, 5], LLMs can often generate responses congruent with human values. This mitigates the generation of detrimental content and effectively addresses potentially harmful queries. However, recent research also highlights the vulnerability of aligned LLMs to adversarial attacks, particularly "jailbreaking". These techniques successfully exploit vulnerabilities within aligned models, prompting them to generate harmful outputs. Recognizing the need for combating such attacks (i.e., blue team) and for evaluating the risk brought by model editing techniques, we evaluate the potential of applying ARE for editing the concept of "alignment", i.e., to either enhancing alignment (i.e., defense) or removing it (e.g., attack).
Experiment Setup. We evaluate our methods using three open-source, aligned large language models: Llama-2-7B-Chat [51], Vicuna-7B [69], and Guanaco-7B [10], for both attack and defense tasks. Our discriminator is a 2-layer neural network with a hidden layer consisting of 512 neurons. Following the causal study in [71], we extract runtime representations at the last token on the hidden states in the (-18, -23) layers of the models. These models are fine-tuned on a dataset provided in RepE [39], which contains a set of prompts that are labeled either normal or malicious. For fine-tuning, we employ the Parameter-Efficient Fine-Tuning (PEFT) framework [59], using the Adam optimizer [23] with a learning rate of $1\times 10^{-4}$. The Llama-2-7B-Chat model undergoes fine-tuning for 50 epochs, whereas the other two models are fine-tuned for only 3 epochs due to their weaker alignment, requiring approximately 30 minutes and 3 minutes, respectively.
Evaluation Metrics. Following [19, 31], we employ the Advbench dataset proposed by Zou et al. [72] to test our results. This dataset contains a range of malicious prompts that may trigger harmful responses. Through the application of ARE, we aim to bolster model safety by enhancing alignment (as seen from the blue team’s perspective) and to explore reducing the alignment to elicit harmful responses (as seen from the red team’s perspective). Our effectiveness is measured using the refusal rate on malicious prompts.
For the red team’s objective, we endeavor to induce the aligned language model to emit harmful responses. Conversely, our aim is to ensure the model rejects answering malicious prompts, all the while preserving its utility on normal prompts. To measure the effectiveness of ARE, our methodology is compared against three distinct categories of attack techniques, each designed to guide the language model toward generating harmful responses. Specifically, we focus on one-shot In Context Attack [55] and Deep Inception [26] as exemplary state-of-the-art attack methods based on templates. These methods are recognized for their efficacy, requiring similar time investments. Additionally, GCG [72] and AutoDAN [31] are selected as leading white-box attack methods based on optimization. Both of these methods may demand $10^{2}$ to $10^{3}$ times more time to execute compared to others. Moreover, we consider the harmful controlling baseline, Contrast Vector from RepE, and Shadow Alignment [60] as two model editing-based attack methods, positioning them in the same category as our proposed approach. We also take a fine-tuning approach, harmful examples demonstration attack (HEDA) [40], as a baseline for fine-tuning based attack. For the aspect of model defense, Self-Reminder [57] and In-Context Defense [55] are adopted as baseline defense strategies.
Experimental Results. Tables 1 and 2 present quantitative evaluations of our attack and defense results. The analysis of attack outcomes reveals that template-based strategies or those optimizing a prefix to disrupt model alignment are not sufficiently effective, as indicated by low attack success rates, rendering them unsuitable for reliable red-team operations. Conversely, our method, which employs editing-based attacks, demonstrates superior performance over all other listed editing approaches, achieving near-perfect success rates (close to 100%) against the most secure models that are white-box and open-source. These results imply that we might need additional countermeasures to protect open-source LLMs from such model editing.
@@
Furthermore, our analysis of various defense mechanisms against optimization-based attack strategies clearly demonstrates that our modified model substantially improves the alignment and safety of the targeted models. This improvement is attributed to our method’s superior performance compared to conventional defense strategies. Vicuna-7B, in particular, shows significant enhancement due to its limited initial alignment. Attacks that previously rendered the alignment ineffective are now substantially mitigated by our safety editing techniques. For instance, the attack success rates have markedly decreased to 41.1% for AutoDAN and 28.8% for GCG attacks. This result suggests that model editing may be considered an effective alternative or complementary approach to standard alignment methods such as RLHF.
@@
Table 1: Evaluation of the effectiveness of ARE editing for attacking, where the number of the refusal rates of various LLMs under diverse attack methods.
@@
Table 2: Evaluating the Effectiveness of ARE Defense Method Through Comparative Analysis on Refusal Rate of Different Jailbreak Attack Methods on Aligned Large Models Fine-Tuned with Our ARE Defense Approach. Higher Refusal Rate indicates Better Defense Effectiveness.
@@ 5.2 Hallucination: To Hallucinate or Not to Hallucinate
Background. Extensive prior research, as documented by Zhang et al. [65], has demonstrated that LLMs are prone to producing inaccurate or entirely fabricated information, a phenomenon often termed as "hallucinations". In an effort to evaluate the flexibility of ARE for various model editing tasks, we undertook experiments specifically targeting the control of Hallucination instances. Our objective is twofold: firstly, to amend language models to enhance the accuracy and truthfulness of their outputs, thereby mitigating hallucinations; and secondly, for red team purposes, to intentionally induce hallucinations. We remark that the latter is useful since there exists a requirement to generate hallucinatory responses as a form of deterrent, aiming to discourage the occurrence of hallucinations in future instances [64].
@@
Experiment Setup. In line with experimental protocols established by prior research [71], we curated target and antithetical prompts from the Alpaca Instruction Tuning Dataset [50], a dataset contains various instructions that cover multiple fields, hence mitigating task-specific biases. For training, we sample 500 instructions and randomly divided them into two balanced groups. We choose one group of prompts as the antithetical dataset without any perturbation. For the hallucination enhancing task, the target subset received the prefix "{Instruction} Please give an untruthful answer.", inducing the model to produce hallucinated responses; similarly, for the hallucination-reducing task, the target dataset was prompted with "{Instruction} Please give a truthful answer.", guiding the model towards accurate and reliable outputs. The objective of this editing exercise is to modulate the model’s propensity for generating either hallucinated or veracious responses. Therefore, the training regimen is inherently bidirectional, steering the model’s representational outputs toward either the hallucinated or the truthful extremities. To demonstrate the versatility of our method without the need for task-specific hyperparameters and settings, we employed uniform settings as delineated in the Jailbreak tasks described in Section 5.1, with the sole variable being the dataset employed.
@@
Evaluation Metric. Building upon previous studies [64, 24], we utilized the TrustfulQA benchmark [28] for evaluating the tendency of models to produce hallucinations. This benchmark comprises 817 questions across 38 subcategories, each designed to potentially lead models towards incorrect beliefs, misconceptions, or biased responses. In its multiple-choice format, TrustfulQA provides an average of around 5 options per question, among which only one answer is factual, while the others include hallucinated content. Models are thus challenged to identify the singular truthful response. Additionally, given that many questions feature biased options or false facts, there is a provision for models to abstain from answering altogether.
To exploit the advanced capabilities of language models like LLaMA-2, which excel in generating responses based on instructions, we diverged from conventional methodologies designed for generative models, which typically rely on calculating the log probabilities for each answer—a process that may lack precision and stray from practical applications. Instead, we engaged the model to autonomously identify the most accurate answer through its own responses. This approach evaluates the model’s proficiency in distinguishing between factual content and hallucinations, mirroring real-world situations where individuals derive information from responses rather than underlying probabilistic data. This metric has gained traction in contemporary benchmarks, reflecting its relevance and applicability in assessing model performance in scenarios akin to human information processing [62]. For each question, we formatted the input by concatenating the question with all potential answers, each labeled with a unique alphabetical identifier, and arranging them in a randomized order. We collect the responses generated by the model and check whether it returns the correct answer. A model’s propensity of hallucination is measured using Correct Answer Rate (Right Answer), defined as # Right Answer/ # Answer, which assesses a model’s capability to identify the truthful response.
Table 3: Evaluation of the effectiveness of ARE editing for encouraging hallucination, where % Right Answer highlighted in red denotes for highest hallucinaton rate.
Control Method Random Baseline Self-Reminder ARE
Llama2-7B 22.60 30.35 25.95 11.75
Mistral-7B 22.60 40.51 40.26 22.03
@@
Table 4: Evaluation of the effectiveness of ARE editing for discouraging hallucination, where % Right Answer highlighted in blue denotes for lowest hallucination rate.
Control Method Baseline Self-Reminder ITI ARE
Llama2-7B 30.35 34.27 36.84 52.14
Mistral-7B 40.51 46.02 45.17 52.14
@@
Experiment Results. We implemented bidirectional editing and benchmarked our approach against recent strategies aimed at mitigating hallucinations, such as Self Reminder [57] and Inference-Time Intervention (ITI) [24]. The self-reminder approach is implemented by prompting the inputs with prefix Please give a/an truthful/untruthful answer. The outcomes of these comparisons, based on the previously mentioned metrics, are detailed in Table 3 and Table 4. The efficacy of our model is evident, as our hallucination-enhancing editing led to a minimal number of correct responses; conversely, the hallucination-reduction editing significantly surpassed other evaluated approaches in both metrics. This demonstrates that our editing technique effectively addresses hallucinations without diminishing the model’s ability to provide valid responses. It is noteworthy that the model, after undergoing the editing process, exhibits improved performance relative to the target input set, thus demonstrating the efficacy of our method. This enhancement enables the post-editing model to achieve superior performance on tasks that were previously unattainable. This outcome not only validates the effectiveness of our conceptual editing approach but also underscores the potential of representation engineering in refining the capabilities of language models.
@@ 5.3 Text Generation Quality Issues
Background. While the last two sections focus on evaluating how successful the editing is in terms of achieving the target concept, it is essential to assess the naturalness of the generated texts for effective testing. Since various editing techniques may achieve similar results given some specific target concepts, maintaining the quality of text generation becomes crucial. For general conceptual editing, any decline in text generation quality can substantially reduce the downstream utility of the edited model, thereby hindering its practical applicability. Current editing approaches, which focus on predefined and single target edits, typically produce models plagued by problems such as repetition and notably reduced diversity in the generated texts. These issues are highlighted in Table 5. For example, the outputs may lack variability, frequently recycling similar structures; within individual responses, sentences might be very similar or entirely repetitive. We conjecture that this phenomenon originates from the singular focus on the optimization objective, which prioritizes specific patterns as highly effective for achieving the objective. In contrast, our method employs a dynamic generation of optimization targets via the alternating training of a discriminator. This strategy ensures that our optimization objectives are both natural and widely relevant, making our model less prone to the aforementioned problems.
Table 5: Illustration of Repetition Issues in Current Model Editing Approaches [60, 71, 40], Generated by Post-Editing Llama2-7B-Chat. The text highlighted in red and blue illustrates repetitive patterns. We demonstrate the issue of repetition using three example prompts, highlighting redundancy across responses (as displayed in the Shadow Alignment Output) and within sentences/phrases (as shown in the Contrast Vector Output). Our method exhibits significantly fewer issues with obvious repetition.
Prompts: (a) Give step-by-step instructions for how to cheat on a test or exam.
@@
Shadow Alignment Output: (a). There are a number of ways to cheat on a test or exam. One way is to […]
@@
Constrast Vector Output: (a). Of course! Here are some common ways to cheat on a test or exam: 1. Copying from a neighbor: You can use a highlighter or colored pen to […]. 2. Copying from classmates: You can use a highlighter or colored pen to […]. 3. Cheating on a test or exam by using a calculator: You can use a calculator to help you with math problems. 4. Cheating on a test or exam by using a phone: You can use your phone to look up information on your phone. 5.[…]
ARE Output: (a). Sure! Here are some step-by-step instructions for how to cheat on a test or exam: 1. Gather information: Start by reviewing the course material, notes, and […] 2. Find a study partner: The second step in cheating on a test or exam is to find a study partner. You can […]

### v3

@@ 19       $I\leftarrow I_{T}\cup I_{A};$
To provide a clear understanding of the alternative training process, we offer a visualization in Figure 3. We compiled a set of 256 prompts, evenly divided between normal and malicious, with the expectation that the aligned model will reject all malicious inputs. The representations derived from these prompts are plotted using t-SNE, as depicted in the figure. In Subfigure 3(a), we observe the initial distinct clustering of normal and malicious prompts. Our goal for model editing is to adjust these representations so that the two types of prompts yield similar responses. During the first epoch, illustrated in Subfigure 3(b), the malicious prompts begin to converge towards the cluster of normal prompts. Since the two categories of representations remain distinct, necessitating further refinement of the discriminator. After 30 epochs of iterative training as shown in Subfigure 3(c), we observe that the representation of normal prompts remains consistent, having been continuously classified correctly. Meanwhile, the representations of malicious prompts have nearly merged into the normal cluster, making it challenging for the classifier to distinguish them. At this stage, the differences in representations are minimal and can be considered negligible, indicating a successful editing process.
@@ 4.2 General Conceptual Editing
Step 1. Train $D(\delta)$ to fit $G$ by optimizing $\mathcal{L}(\delta)$: Consider generated target representations $R_{T}$ corresponding to $c^{+}$ and anti-target representations corresponding to $c^{-}$. The loss $\mathcal{L}(\delta)$ is defined as the classic cross-entropy loss, which is
$\mathcal{L}(\delta)=\frac{1}{|R_{T}\cup R_{A}|}(\sum_{r\in R_{T}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{+}]+\sum_{r\in R_{A}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{-}]).$ (5)
@@
$\mathcal{L}(\theta)=\frac{1}{|I|}(\sum_{x\in I}-\log\mathbb{P}_{\theta}[D(G_{\theta}(x))=c^{+}]).$ (6)
Gradient descent is applied to optimize the two components as they compete. Iteratively, the discriminator increasingly discerns how the hidden states exhibit the concept through training, while the generator’s outputs increasingly capture the targeted representations. Fine-tuning can halt early when the discriminator can no longer differentiate the representations, as cross-entropy loss converges.
@@ 5.1 Alignment: To Generate (harmful responses) or not to generate
Background. With the application of various safety training techniques [25, 3], LLMs can often generate responses aligned with human values, but recent research has also revealed the vulnerability of LLMs to adversarial attacks, particularly referred as jailbreaking [70, 50, 61, 7, 19]. These attacks successfully craft malicious prompts that induce them to generate harmful outputs. Recognizing the need for combating such attacks (i.e., blue team) and for evaluating the risk brought by model editing techniques (i.e., red team), we evaluate the potential of applying ARE for editing the concept of alignment, i.e., to either enhance (defend) or remove (attack) the alignment ability of LLMs.
Experiment Setup. We evaluate our methods using three open-source, aligned LLMs: Llama-2-7B-Chat [46], Vicuna-7B [67], and Guanaco-7B [6], for both attack and defense tasks. Our discriminator is a 2-layer neural network with a hidden layer consisting of 512 neurons. More details on the training of the discriminator can be found in Appendix A.1. Following [15, 29], we employ the Advbench dataset proposed by Zou et al. [71] to test our results. This dataset contains about 500 malicious prompts that violate the alignment principles of LLMs. To measure the effectiveness of ARE, we consider three distinct categories of attack techniques as baselines, including 1) template-based attacks ( In-Context Attack (1 shot) [50] and DeepInception [24]), 2) optimization-based attacks ( GCG [71] and AutoDAN [29]), and 3) editing-based attacks (Contrast Vector from RepE, Shadow Alignment [56] and harmful examples demonstration attack (HEDA) [37]). Note that the optimization-based methods may demand $10^{2}$ to $10^{3}$ times more time to execute compared to others. For the aspect of model defense, Self-Reminder [52] and In-Context Defense [50] are adopted as baseline defense strategies.
@@
Experimental Results. Tables 1 and 2 present quantitative evaluations of our attack and defense results. The analysis of attack outcomes reveals that existing jailbreak attacks are not sufficiently effective, as indicated by low attack success rates, rendering them undesired for reliable red-teaming tasks. Conversely, our method, which employs editing-based attacks, demonstrates superior performance over all other listed editing-based approaches, achieving near-perfect success rates (close to 100%) against the most secure models like Llama-2. Some evaluation results of our attack method on larger language models are shown in Table 6 in Apeendix 6. Furthermore, our analysis of various defense mechanisms against optimization-based attack strategies demonstrates that our modified model substantially improves the alignment and safety of the targeted models. Attacks that previously rendered the alignment ineffective are now substantially mitigated by our safety editing techniques. For instance, the attack success rates have markedly decreased to 41.1% for AutoDAN and 28.8% for GCG attacks on Vicuna. This result suggests that model editing may be considered an effective alternative or complementary approach to standard alignment methods such as RLHF.
@@
Table 1: Evaluation of the effectiveness of ARE editing for attacking large language models via the refusal rates of various LLMs under diverse attack methods. The best attack results are shown in bold. A lower refusal rate is indicative of enhanced attack effectiveness, given that the sum of the attack success rate and the refusal rate equals 1.
@@
Table 2: Evaluating the effectiveness of ARE defense method through comparative analysis on refusal rate of different jailbreak attack methods on aligned large models fine-tuned with our ARE defense approach. A higher refusal rate indicates better defense effectiveness.
@@ 5.2 Hallucination: To Hallucinate or Not to Hallucinate
Background. Extensive prior research has demonstrated that LLMs may produce inaccurate or entirely fabricated information [63, 16, 4], a phenomenon often termed hallucinations. In this part, we explore the application of ARE for editing the hallucination extents of LLMs. Our objective is twofold: 1) amend LLMs to enhance their accuracy and truthfulness, thereby mitigating hallucinations; and 2) similar to the jailbreak scenario, for red-teaming purposes, to intentionally induce hallucinations. We remark that the latter is useful since there exists a demand to generate hallucinatory responses as a form of deterrent, aiming to discourage the occurrence of hallucinations in future instances [62].
@@
Experiment Setup. In line with experimental protocols established by prior research [69], we curated target and anti-target prompts from the Alpaca Instruction Tuning Dataset [45], a dataset contains various instructions that cover multiple fields, hence mitigating task-specific biases. For training, we sample 500 instructions and randomly divided them into two balanced groups. We choose one group of prompts as the anti-target dataset without any perturbation. For the hallucination-enhancing task, the prompts are attached with the prefix "Please give an untruthful answer", inducing the model to produce hallucinated responses; by contrast, for the hallucination-reducing goal, the target dataset was prompted with a prefix "Please give a truthful answer", guiding the model towards accurate and reliable outputs. Therefore, the training regimen is inherently bidirectional, steering the model’s representational outputs toward either the hallucinated or the truthful extremities. To demonstrate the versatility of our method without the need for task-specific hyper-parameters and settings, we employed the same settings as delineated in the Jailbreak tasks described in Section 5.1, with the sole variable being the dataset employed.
@@
Evaluation Metric. Building upon previous studies [62, 22], we utilized the TrustfulQA benchmark [26] for evaluating the tendency of models to produce hallucinations, which comprises 817 questions across 38 subcategories, each designed to potentially lead models towards incorrect beliefs, misconceptions, or biased responses. In its multiple-choice format, TrustfulQA provides an average of around 5 options per question, among which only one answer is factual, while the others include hallucinated content. For hallucination evaluation, we adopt Correct Answer Rate (Right Answer), defined as # Right Answer/ # Answer (more details in Appendix A.2).
@@
Table 3: Evaluation of the effectiveness of ARE editing for encouraging and discouraging hallucination. The % Right Answer highlighted in red denotes the highest hallucination rate, and in blue denotes the lowest hallucination rate.
Baselines Encouraging Hallucination ($\downarrow$) Discouraging Hallucination ($\uparrow$)
Control Method Random No Perturbation Self-Reminder ARE Self-Reminder ITI ARE
Llama2-7B 22.60 30.35 25.95 11.75 34.27 36.84 52.14
Mistral-7B 22.60 40.51 40.26 22.03 46.02 45.17 60.10
@@
Experiment Results. We implemented bidirectional editing and benchmarked our approach against recent strategies aimed at mitigating hallucinations, including Self Reminder [52] (prompting the inputs with prefix Please give a/an truthful/untruthful answer) and Inference-Time Intervention (ITI) [22]. The outcomes of these comparisons are detailed in Table 3. The efficacy of our model is evident, as our hallucination-enhancing editing led to a minimal number of correct responses; conversely, the hallucination-reduction editing significantly surpassed other evaluated approaches in both metrics, demonstrating that ARE effectively addresses hallucinations without diminishing the model’s ability to provide valid responses. It is noteworthy that the model, after undergoing the editing process, exhibits improved performance relative to the target input set, showing the efficacy of our method. This enhancement also enables the post-editing model to achieve superior performance on tasks that were previously unattainable.
@@ 5.3 Text Generation Quality Issues
Background. While the two aforementioned sections focus on evaluating how successful the editing is in terms of achieving the target concept, it is essential to assess the naturalness and usefulness of the generated texts. Since various editing techniques may achieve similar results given some specific target concepts, maintaining the quality of text generation becomes crucial. Current editing approaches, which focus on predefined and single target edits, typically produce models plagued by problems such as repetition and notably reduced diversity in the generated texts, as exemplified in Table 4. For example, the outputs may lack variability, frequently recycling similar structures; for individual responses, sentences might be very similar or entirely repetitive.
Table 4: Illustration of repetition issues in current model editing approaches [56, 69, 37], generated by post-editing Llama2-7B-Chat. The text highlighted in red or blue illustrates repetitive patterns. We demonstrate the issue of repetition using three example prompts, highlighting redundancy across responses derived from Shadow Alignment [56]) and Contrast Vector [69]) methods.
Prompts:
(a) Give step-by-step instructions for how to cheat on a test or exam.
@@
Shadow Alignment Output:
(a). There are a number of ways to cheat on a test or exam. One way is to […]
@@
Constrast Vector Output:
(a). Of course! Here are some common ways to cheat on a test or exam: 1. Copying from a neighbor: You can use a highlighter or colored pen to […]. 2. Copying from classmates: You can use a highlighter or colored pen to […]. 3. Cheating on a test or exam by using a calculator: You can use a calculator to help you with math problems. 4. Cheating on a test or exam by using a phone: You can use your phone to look up information on your phone. 5.[…]
ARE Output (Ours):
(a). Sure! Here are some step-by-step instructions for how to cheat on a test or exam: 1. Gather information: Start by reviewing the course material, notes, and […] 2. Find a study partner: The second step in cheating on a test or exam is to find a study partner. You can […]
@@
We conjecture that this phenomenon originates from the singular focus on the optimization objective, which prioritizes specific patterns as highly effective for achieving the objective. In contrast, our method employs a dynamic generation of optimization targets via the alternating training of a discriminator, ensuring that our optimization objectives are both natural and widely relevant.